OpenStudy (anonymous):

What are the x- and y-intercepts of the quadratic equation y = (x - 3)2 - 25? Identify the x-intercepts by graphing the quadratic equation

OpenStudy (anonymous):

Simplify the equation to get y=2x-31.Rearrange to get 2x-y=31.Now divide on both sides by 31 to get 2x/31+(-y/31)=1.Simplify more to x/(31/2)+y/(-31).Now compare it to x/a+y/b=1 which is the intercept form to get x intercept as 31/2 and y intercept as -31

OpenStudy (anonymous):

y=0 to find x-intercepts this is difference of 2 squares (x-3-5)(x-3+5)=0 (x-8)(x+2)=0 x=8,-2

OpenStudy (anonymous):

@Dockworker you put y=0 you then get (x-3)2-25=0 which means x=31/2

OpenStudy (anonymous):

@pavankumartgpk the 2 in this equation is exponent

OpenStudy (anonymous):

Oh I see.